# Introducing Packet-Level Analysis in Programmable Data Planes to Advance Network Intrusion Detection

### v1

@@ Abstract
Programmable data planes offer precise control over the low-level processing steps applied to network packets, serving as a valuable tool for analysing malicious flows in the field of intrusion detection. Albeit with limitations on physical resources and capabilities, they allow for the efficient extraction of detailed traffic information, which can then be utilised by Machine Learning (ML) algorithms responsible for identifying security threats. In addressing resource constraints, existing solutions in the literature rely on compressing network data through the collection of statistical traffic features in the data plane. While this compression saves memory resources in switches and minimises the burden on the control channel between the data and the control plane, it also results in a loss of information available to the Network Intrusion Detection System (NIDS), limiting access to packet payload, categorical features, and the semantic understanding of network communications, such as the behaviour of packets within traffic flows. This paper proposes P4DDLe, a framework that exploits the flexibility of P4-based programmable data planes for packet-level feature extraction and pre-processing. P4-empowered Ddos detection with Deep Learning (P4DDLe) leverages the programmable data plane to extract raw packet features from the network traffic, categorical features included, and to organise them in a way that the semantics of traffic flows is preserved. To minimise memory and control channel overheads, P4DDLe selectively processes and filters packet-level data, so that all and only the relevant features required by the NIDS are collected. The experimental evaluation with recent Distributed Denial of Service (DDoS) attack data demonstrates that the proposed approach is very efficient in collecting compact and high-quality representations of network flows, ensuring precise detection of DDoS attacks.
@@
Programmable Data Planes, Network Intrusion Detection, Packet-Level features, Machine Learning, DDoS attacks
@@ 1 Introduction
This concept has been extensively explored in the scientific literature, with several studies proposing solutions based on the aggregation of traffic statistics computed within the data plane. The primary objective of these solutions is to optimise the load on the control channel and effectively manage memory utilization in the data plane [17, 18, 19, 20, 21]. One notable drawback of relying solely on statistical traffic features, such as flow duration, averages, maximums, minimums, and standard deviations of attributes like packet size, rate, and inter-arrival time, is the inability to report categorical features (e.g., IP and TCP flags) or portion of packets’ payload directly to the control plane. These packet-level traffic features are crucial for certain ML-based NIDSs in the current state of the art (e.g., [22, 23, 24, 25, 26], among many others), which rely on such traffic attributes to learn the semantic of malicious traffic flows and to segregate them from legitimate network activities.
@@
The remainder of the paper is structured as follows. Section 2 reviews and discusses the related work. Section 3 provides an overview of P4 data plane programming, counting Bloom filters and the CNN used in our experiments. Section 4 presents P4DDLe’s architecture and the methods for feature extraction and storage. Section 5 details the setup of the experimental evaluation presented in Section 6. Finally, the conclusions are given in 7.
@@ 3.1 Data Plane Programming with P4
[…]
Figure 2: P4-programmable pipeline components.
As shown in the figure, the Match-Action block may consist of one or multiple Match-Action Units. Each MAU comprises TCAM and SRAM memory blocks (MEM in the figure) for various purposes such as match-action tables, registers, meters, etc. The Arithmetic Logic Units blocks instead execute arithmetic operations and modifications on the packets’ headers and metadata based on the content of the match-action tables. ALUs may use stateful objects stored in the SRAM memory (e.g., registers) for additional operations.
Relevant to this work are the stateful objects of type register. Registers belong to a wider category of stateful elements called extern objects, which also include counters and meters. Unlike match-action tables, which can be modified only by the control plane, extern objects can be read and written from both data and control planes. Therefore, the access to registers, and to extern objects in general, should be managed with atomic operations to prevent race conditions.
@@ 4 System Architecture
Given these premises, we present P4DDLe, an approach to intrusion detection for P4-based programmable networks that enables resource-efficient traffic feature extraction and filtering in the data plane. The key idea of P4DDLe is to extract and store in the data plane only the features that are essential to the Traffic Classifier running on the control plane, thereby optimising the amount of data transmitted through the control channel. For this purpose, we have designed a system architecture matching the PISA model presented in Section 3.1, with the support of control plane software for traffic classification (Figure 3). The P4 program executed in the data plane extracts the packet attributes and makes them available to the control plane. On the other hand, the control plane is responsible for gathering the traffic metadata from the registers allocated in the data plane and coordinating the read/write access to such registers.
@@ 4.4 Control-Data Plane interaction
Figure 5 illustrates the timelines of Control and Data Plane operations. The two planes operate in parallel and never block each other. The Data Plane always writes on the same block until the value of the Switch Register is changed from the Control Plane. If we start from time $t_{0}$, as shown in the figure, once the control plane sets the value of the Switch Register to $1$, the Data Plane will start writing on memory block $B_{1}$. After that, the Control Plane will first read block $B_{0}$ and, once done with reading the block, it will reset all the registers of blocks $B_{0}$, $F_{0}$ and $c_{0}$ and the traffic features will be processed for classification. The whole process restarts at time $t_{3}$, when the Control Plane sets the Switch Register to 00 to retrieve the features collected in memory block $B_{1}$, while the Data Plane switches to block $B_{0}$.
@@ 6 Evaluation Results
One of the key benefits of P4DDLe over other approaches is the ability to extract raw packet features from the network traffic, categorical features included, and to organise them in a way that the semantic of traffic flows is preserved. P4DDLe efficiently achieves this objective by implementing a counting Bloom filter (Section 3.2) that tracks the number of packets/flow without any wastage of memory resources in the data plane. We demonstrate the advantages of P4DDLe through simulation and emulation experiments, where we disable the tracking methods and observe the corresponding changes using a range of metrics. In the simulation scenario, no actual network data is involved. Instead, we rely on the network profile of a campus network [60] and we demonstrate the benefits of P4DDLe by simulating the feature extraction and storage process in the data plane. The second set of experiments has been conducted in an emulated environment consisting of two hosts (an attacker and a victim) and a P4-enabled software switch with configurable registers. The experiments involve injecting pre-recorded network traffic into this environment, using a publicly available dataset of both benign and DDoS traffic.
@@ 6.1 Simulation Scenario
We utilise the profile outlined in [60] to simulate the traffic of a realistic network. This profile was generated using network activity data gathered from a university campus. From it, we extracted the distribution of TCP, UDP, and other protocols and we computed the average number of packets/flow for each protocol.
Each experiment consists of $1000$ iterations, each simulating the collection of a random number of flows, ranging from 1 to 8192 flows. The length of such flows (number of packets/flow) is generated based on the aforementioned traffic profiles, while the arrival of packets across different flows has been randomised to ensure a non-sequential packet distribution. To minimise potential issues due to the insufficient space in memory to store packets for all the 8192 flows (which is not the goal of this analysis), we reserve space in the memory block $B_{k}$ for at least two packets/flow by setting the value $n=16384$ packets (see also Figure 4 for reference).
We execute the experiments with P4DDLe and we compare the results against those obtained using a baseline configuration, in which we disable the P4DDLe’s algorithm for filtering the packets (Algorithm 1).
@@ 6.1.1 Bloom filter size
Each Bloom filter is characterised by two key dimensions: the number of cells, denoted as $m\in[0,2^{r}-1]$, and the size of each cell. In this experiment, we have set the size of the cells while varying the value of $r$ to understand the sensitivity of P4DDLe to the number of cells in terms of Collected Flows. Considering that 75% of the flows within the traffic profile consist of no more than 4 packets, we have chosen to set the cell size to 3 bits. This size adequately accommodates the counting of up to 4 packets/flow. Nevertheless, in the next section, we evaluate the sensitivity of P4DDLe to variations in the number of packets/flow and the cell size.
@@
Figure 7 presents the results obtained with Bloom filter sizes from $1024$ to $32768$ cells (or $r$ from $10$ to $15$). As expected, the larger the Bloom filter size, the smaller the number of hash collisions and the higher the number of flows collected. Remarkably, it is worth noting that P4DDLe is able to collect more flows than the baseline, regardless of the Bloom filter size. It is important to recall that the baseline approach gathers all packets without any filtering logic. Considering that the average number of packets/flow in the traffic profile is 78, the baseline approach exhausts the memory capacity with approximately $210$ flows (calculated as $16384/78$). Consequently, once the memory block reaches its limit, older packets are overwritten by more recent ones, resulting in a loss of flows.
@@ 6.1.2 Maximum number of collected packets per flow
In this experiment, we fix the size of the Bloom filter to $32768$ cells, and we vary $p$ from $2$ to $32$ packets/flow. We expect that by decreasing the value of $p$, the average number of packets collected per flow will also decrease, leading to a higher number of collected flows in memory but also to more potential collisions.
@@ 6.2 Emulated Scenario
In this experiment, we generate a series of DDoS attacks by using the traffic traces of the CIC-DDoS2019 dataset (see Section 5.1). If we exclude the DNS attack, the flows of the other attacks present an average flow length of around $2.2$ packets, hence considerably smaller than those of the profile of the campus network used in simulations (around $78$ packets/flow). Considering that we chose $p=4$ (i.e., $p>2.2$), we expect that with these small flows, both P4DDLe and the baseline can collect approximately the same number of flows. Acknowledging that malicious traffic flows typically do not constitute the totality of network traffic, we define four evaluation scenarios to enable a comprehensive comparison. These scenarios involve the inclusion of background benign traffic at varying proportions relative to the total packets in the network, namely 0% (no benign traffic), 25%, 50%, and 75%. Table II presents the average flow length as we increase the percentage of benign traffic in the network (whose average flow length is about $36$ packets).
@@
Other important information pertains to the relationship between packets and flows. Since we add a portion of background benign traffic based on the total amount of packets, this does not necessarily reflects to an analogous increase in the number of flows. Actually, based on the low average flow length in the attacks (first column in Table II), and the relatively high number of packets/flow of the benign trace (around $36$ on average), even splitting 50% between benign and DDoS packets, the average percentage of DDoS flows is higher than 93% on the total of flows (excluding the DNS attack that presents an exceptional behaviour). Table III shows the percentage of DDoS flows for each attack type.
@@ 6.2.1 Collected Flows and sFNR
Figure 9 shows the comparison related to Collected Flows for all attack types. First, we can notice that with the baseline approach the percentage of collected flows remains stable when we vary the proportion between benign and attack traffic. This can be attributed to the absence of any filtering logic, causing the percentage of collected flows to solely depend on the memory capacity. In contrast, P4DDLe demonstrates consistent improvement in the percentage of collected flows as the volume of benign traffic (and the average flow length) increases. While the baseline approach fills up the memory block predominantly with large benign flows, P4DDLe optimizes memory usage by storing only the essential information required by the NIDS and filtering out packets that exceed the maximum limit of $p=4$. This memory management prevents large flows from occupying excessive memory space, thus ensuring that a higher number of flows can be transmitted to the NIDS for classification.
For instance, in the case of the MSSQL attack, whose average flow size is around $2$ packets/flow, with 0% of benign traffic the performance of P4DDLe is approximately the same as with the baseline approach. This is because with $p=4$, no packets are filtered and the memory occupation is similar with both approaches. However, as soon as we add some benign traffic, which is composed of larger flows, the filtering mechanism starts dropping the extra packets, saving space in memory for more flows. This is reflected in the sFNR metric (Figure 10), which shows the ability of P4DDLe to collect more malicious flows than the baseline approach, ultimately leading to faster mitigation of network attacks.
It is also worth noticing that when the volume of the attack is small, like in the case of LDAP and DNS, the memory is never filled, even when adding benign traffic. Therefore, the collected flows is 100% for both, while the non-zero sFNR obtained with the DNS attacks is solely due to Lucid’s FNR.

### v4

@@ Abstract
Programmable data planes offer precise control over the low-level processing steps applied to network packets, serving as a valuable tool for analysing malicious flows in the field of intrusion detection. Albeit with limitations on physical resources and capabilities, they allow for the efficient extraction of detailed traffic information, which can then be utilised by Machine Learning (ML) algorithms responsible for identifying security threats. In addressing resource constraints, existing solutions in the literature rely on compressing network data through the collection of statistical traffic features in the data plane. While this compression saves memory resources in switches and minimises the burden on the control channel between the data and the control plane, it also results in a loss of information available to the Network Intrusion Detection System (NIDS), limiting access to packet payload, categorical features, and the semantic understanding of network communications, such as the behaviour of packets within traffic flows. This paper proposes P4DDLe, a framework that exploits the flexibility of P4-based programmable data planes for packet-level feature extraction and pre-processing. P4-empowered Ddos detection with Deep Learning (P4DDLe) leverages the programmable data plane to extract raw packet features from the network traffic, categorical features included, and to organise them in a way that the semantics of traffic flows are preserved. To minimise memory and control channel overheads, P4DDLe selectively processes and filters packet-level data, so that only the features required by the NIDS are collected. The experimental evaluation with recent Distributed Denial of Service (DDoS) attack data demonstrates that the proposed approach is very efficient in collecting compact and high-quality representations of network flows, ensuring precise detection of DDoS attacks.
@@
Programmable Data Planes, Network Intrusion Detection, Packet-Level features
@@ 1 Introduction
This concept has been extensively explored in the scientific literature, with several studies proposing solutions based on the aggregation of traffic statistics computed within the data plane. The primary objective of these solutions is to optimise the load on the control channel and effectively manage memory utilization in the data plane [17, 18, 19, 20, 21]. One notable drawback of relying solely on statistical traffic features, such as flow duration, averages, maximums, minimums, and standard deviations of attributes like packet size, rate, and inter-arrival time, is the inability to report categorical features (e.g., IP and TCP flags, Differentiated Services Code Point (DSCP), etc.), the timestamp of packets or portion of packets’ payload directly to the control plane. These packet-level traffic features are crucial to detect some types of network intrusions, such as brute force attacks (e.g., by monitoring the TCP flags, including SYN, ACK, FIN and RST flags [22]), anomalous QoS settings (monitoring the DSCP code points) [23], TCP, UDP and ICMP fragmentation attacks (monitoring the IP flags) [24], etc. In addition, the timestamp of packets and portions of the payload are employed in combination with categorical features by certain ML-based NIDSs in the current state of the art (e.g., [25, 26, 27, 28, 29], among many others) to learn the semantic of malicious traffic flows and to segregate them from legitimate network activities.
@@
The main contributions of this work can be summarised as follows:
  1. 1.
    A framework for P4-based programmable data planes called P4DDLe, designed to cope with the requirements of modern NIDSs, which often rely on raw packet-level features for accurate cyber threat detection.
  2. 2.
    A mechanism based on counting Bloom filters to efficiently group packet-level features into traffic flows while minimizing the chances of collisions between the flow identifiers.
  3. 3.
    A comprehensive assessment of the proposed approach, including a sensitivity analysis of the Bloom filter size (directly linked to the probability of collisions), and the evaluation of a state-of-the-art NIDS in detecting realistic DDoS attacks using P4DDLe’s packet-level representation of the network traffic.
The remainder of the paper is structured as follows. Section 2 reviews and discusses the related work. Section 3 provides an overview of P4 data plane programming, counting Bloom filters and the CNN used in our experiments. Section 4 presents P4DDLe’s architecture and the methods for feature extraction and storage. Section 5 details the setup of the experimental evaluation presented in Section 6. Finally, the conclusions are given in Section 7.
@@ 3.1 Data Plane Programming with P4
P4 is a language for expressing how the network traffic is processed by the data plane of programmable network elements such as hardware and software switches. P4 is target-independent, thus it supports a variety of targets such as ASICs, FPGAs, NICs and software switches.
The basic components of a P4-programmable forwarding pipeline are illustrated in Figure 2. This diagram outlines the Version 1.0 Switch Architecture Model [48], abbreviated as V1model, which is widely adopted by developers for P4 program implementation. We utilised this model as a reference for P4DDLe. The pipeline encompasses a Parser, a state machine that extracts packet headers and metadata from the incoming bitstream, a Checksum verification control block, an Ingress Match-Action processing control block, an Egress Match-Action processing control block, a Checksum update control block and a Deparser, which assembles the headers with the payload received from the Parser.
Figure 2: P4-programmable pipeline components (V1model).
As shown in the figure, the Match-Action block may consist of one or multiple Match-Action Units. Each MAU comprises TCAM and SRAM memory blocks (MEM in the figure) for various purposes such as match-action tables, registers, meters, etc. The Arithmetic Logic Units blocks instead execute arithmetic operations and modifications on the packets’ headers and metadata based on the content of the match-action tables, which host the traffic forwarding rules. ALUs may use stateful objects stored in the SRAM memory (e.g., registers) for additional operations.
Relevant to this work are the stateful objects of type register. Registers belong to a wider category of stateful elements called extern objects, which also include counters and meters. Unlike match-action tables, which are stored in the TCAM memory and can be modified only by the control plane, extern objects in the SRAM memory can be read and written from both data and control planes. Therefore, the access to registers, and to extern objects in general, should be managed with atomic operations to prevent race conditions.
@@ 4 System Architecture
Given these premises, we present P4DDLe, an approach to intrusion detection for P4-based programmable networks that enables resource-efficient traffic feature extraction and filtering in the data plane. The key idea of P4DDLe is to extract and store in the data plane only the features that are essential to the Traffic Classifier running on the control plane, thereby optimising the amount of data transmitted through the control channel. For this purpose, we have designed a system architecture matching the V1model presented in Section 3.1, with the support of control plane software for traffic classification (Figure 3). The P4 program executed in the data plane extracts the packet attributes and makes them available to the control plane. On the other hand, the control plane is responsible for gathering the traffic metadata from the registers allocated in the data plane and coordinating the read/write access to such registers.
@@ 4.4 Control-Data Plane interaction
Figure 5 illustrates the timelines of Control and Data Plane operations. The two planes operate in parallel and never block each other. The Data Plane always writes on the same block until the value of the Switch Register is changed from the Control Plane. If we start from time $t_{0}$, as shown in the figure, once the control plane sets the value of the Switch Register to $1$, the Data Plane will start writing on memory block $B_{1}$. After that, the Control Plane will first read block $B_{0}$ and, once done with reading the block, it will reset all the registers of blocks $B_{0}$, $F_{0}$ and $c_{0}$ and the traffic features will be processed for classification. The whole process restarts at time $t_{3}$, when the Control Plane sets the Switch Register to 00 to retrieve the features collected in memory block $B_{1}$, while the Data Plane switches to block $B_{0}$. It is important to note that the frequency of this process (computed as $1/(t_{3}-t_{0})$ using the example in Figure 5) is managed by the Control Plane and does not depend on Data Plane’s state.
@@ 6 Evaluation Results
One of the key benefits of P4DDLe over other approaches is the ability to extract raw packet features from the network traffic, categorical features included, and to organise them in a way that the semantics of traffic flows are preserved. P4DDLe efficiently achieves this objective by implementing a counting Bloom filter (Section 3.2) that tracks the number of packets/flow without any wastage of memory resources in the data plane. We demonstrate the advantages of P4DDLe through simulation and emulation experiments, where we disable the tracking methods and observe the corresponding changes using a range of metrics.
In the simulation scenario, no actual network data is involved. Instead, we rely on the network profile of a campus network [62] and we demonstrate the benefits of P4DDLe by simulating the feature extraction and storage process in the data plane. The second set of experiments has been conducted in an emulated environment consisting of two hosts (an attacker and a victim) and a P4-enabled software switch with configurable registers. The experiments involve injecting pre-recorded network traffic into this environment, using a publicly available dataset of both benign and DDoS traffic.
@@ 6.1 Simulation Scenario
We utilise the profile outlined in [62] to simulate the benign traffic of a realistic network. This profile was generated using network activity data gathered from a university campus and does not present any documented attack in the trace. From it, we extracted the distribution of TCP, UDP, and other protocols and we computed the average number of packets/flow for each protocol.
Each experiment consists of $1000$ iterations, each simulating the collection of a random number of flows, ranging from 1 to 8192 flows. The length of such flows (number of packets/flow) is generated based on the aforementioned traffic profiles, while the arrival of packets across different flows has been randomised to ensure a non-sequential packet distribution, aligning with the characteristics of realistic packet-switched networks. To minimise potential issues due to the insufficient space in memory to store packets for all the 8192 flows (which is not the goal of this analysis), we reserve space in the memory block $B_{k}$ for at least two packets/flow by setting the value $n=16384$ packets (see also Figure 4 for reference).
These experiments are realised with Python scripts designed to simulate a single collection of network flows. The scripts replicate the logic of both P4DDLe (as described in Algorithm 1) and a baseline configuration. In the baseline setup, the packet filtering algorithm intrinsic to P4DDLe is deliberately deactivated, causing all incoming packets to be stored in memory within the circular buffer, with no constraints on the number of packets per flow. These scripts are publicly available for testing on the P4DDLe repository [54].
@@ 6.1.1 Bloom filter size
Each Bloom filter is characterised by two key dimensions: the number of cells, denoted as $m\in[0,2^{r}-1]$, and the size of each cell. In this experiment, we have set the size of the cells while varying the value of $r$ to understand the sensitivity of P4DDLe to the number of cells in terms of Collected Flows. Considering that 75% of the flows within the traffic profile consist of no more than 4 packets, we have chosen to set the cell size to 3 bits. This size adequately accommodates the counting of up to 4 packets/flow.
@@
Figure 7 presents the results obtained with Bloom filter sizes from $1024$ to $32768$ cells (or $r$ from $10$ to $15$). The figure also includes the results obtained with a single hash function (with $r=15$) , allowing for a comparative analysis with Bloom filters. As expected, the larger the Bloom filter size, the smaller the number of hash collisions and the higher the number of flows collected. Remarkably, it is worth noting that P4DDLe is able to collect more flows than the baseline, regardless of the Bloom filter size. It is important to recall that the baseline approach gathers all packets without any filtering logic. Considering that the average number of packets/flow in the traffic profile is 78, the baseline approach exhausts the memory capacity with approximately $210$ flows (calculated as $16384/78$). Consequently, once the memory block reaches its limit, older packets are overwritten by more recent ones, resulting in a loss of flows. The results of this experiment also confirm the efficiency of Bloom filters over a single hash function in limiting the collisions, as discussed in Section 3.2 and demonstrated by Tarkoma et al. [47].
@@ 6.1.2 Maximum number of collected packets per flow
In this experiment, we fix the size of the Bloom filter to $32768$ cells, and we vary $p$ from $2$ to $128$ packets/flow. We expect that by decreasing the value of $p$, the average number of packets collected per flow will also decrease, leading to a higher number of collected flows in memory but also to more potential collisions.
@@ 6.2 Emulated Scenario
In this experiment, we generate a series of DDoS attacks by using the traffic traces of the CIC-DDoS2019 dataset (see Section 5.1). If we exclude the DNS attack, the flows of the other attacks present an average flow length of around $2.2$ packets, hence considerably smaller than those of the profile of the campus network used in simulations (around $78$ packets/flow). Considering that we chose $p=4$ (i.e., $p>2.2$), we expect that with these small flows, both P4DDLe and the baseline can collect approximately the same number of flows. Acknowledging that malicious traffic flows typically do not constitute the totality of network traffic, we define four evaluation scenarios to enable a comprehensive comparison. These scenarios involve the inclusion of background benign traffic at varying proportions relative to the total packets in the network, namely 0% (no benign traffic), 25%, 50%, and 75%. Table II presents the average flow length as we increase the percentage of benign traffic in the network (whose average packet rate is about 0.7 packets/sec and average flow length is about $36$ packets).
@@
Other important information pertains to the relationship between packets and flows. Since we add a portion of background benign traffic based on the total amount of packets, this does not necessarily reflect an analogous increase in the number of flows. Actually, based on the low average flow length in the attacks (first column in Table II), and the relatively high number of packets/flow of the benign trace (around $36$ on average), even splitting 50% between benign and DDoS packets, the average percentage of DDoS flows is higher than 93% on the total of flows (excluding the DNS attack that presents an exceptional behaviour). Table III shows the percentage of DDoS flows for each attack type.
@@ 6.2.1 Collected Flows and sFNR
Figure 9 shows the comparison related to Collected Flows for all attack types. First, we can notice that with the baseline approach the percentage of collected flows remains stable when we vary the proportion between benign and attack traffic. This can be attributed to the absence of any filtering logic, causing the percentage of collected flows to solely depend on the memory capacity. In contrast, P4DDLe demonstrates consistent improvement in the percentage of collected flows as the volume of benign traffic (and the average flow length) increases. While the baseline approach fills up the memory block predominantly with large benign flows, P4DDLe optimizes memory usage by storing only the essential information required by the NIDS and filtering out packets that exceed the maximum limit of $p=4$. This memory management prevents large flows from occupying excessive memory space, thus ensuring that a higher number of flows can be transmitted to the NIDS for classification. For instance, in the case of the MSSQL attack, whose average flow size is around $2$ packets/flow, with 0% of benign traffic the performance of P4DDLe is approximately the same as with the baseline approach. This is because with $p=4$, no packets are filtered and the memory occupation is similar with both approaches. However, as soon as we add some benign traffic, which is composed of larger flows, the filtering mechanism starts dropping the extra packets, saving space in memory for more flows.
P4DDLe inherently maintains a balanced allocation of memory for both low-rate and high-rate flows, resulting in similar probabilities for their storage. This balance can be observed when analyzing the distribution of collected flows, distinguishing between DDoS traffic, whose packet rate ranges from 2.6K to 30K packets per second, and benign flows, which average around 0.7 packets per second. For instance, consider the SSDP attack scenario with $30$K packets/sec, where the test traffic is divided into $25\%$ benign and $75\%$ attack traffic. In this worst-case scenario, the average percentage of collected flows is approximately $42\%$ for benign flows and $46\%$ for attack flows, demonstrating P4DDLe’s stability and effectiveness in handling heterogeneous network conditions.
The sFNR metric (Figure 10) shows the ability of P4DDLe to collect more malicious flows than the baseline approach, ultimately leading to faster mitigation of network attacks. It is also worth noticing that when the volume of the attack is small, like in the case of LDAP and DNS, the memory is never filled, even when adding benign traffic. Therefore, the collected flows is 100% for both, while the non-zero sFNR obtained with the DNS attacks is solely due to Lucid’s FNR.
6.2.2 Flow quality